Shift of Eve 10-Oct-1999

Mark Timko (timko@kpasa)
Mon, 4 Oct 1999 00:07:13 -0500

Shift Summary for eve shift, 03/Oct/1999

On Shift: Rick Kessler, Mark Timko

Hours of beam available at all intensities: 4.74 hours
Hours of beam available at nominal intensity: 4.75 hours
Hours of E799 data (special runs not included): 4.75 hours
Hours of special runs (laser scans, muon runs...): 3.50 hours

Accounting of lost time:
15 minutes Start new run.
3.5 hours TeVatron down (CHL problems)

Runs ended this shift:
Run 14798 - 13.1M events - BEAM_799
Run 14799 - 9.6M events - BEAM_799
Run 14800 - ........... - Laser Scan

Tapes brought to FCC this shift:
(including tapes for online split)

Accesses during shift:

22:00 Bob T. and Rick F. : into NM2 target to remove lead
shields on accidental counters.

Overview of shift:

Started run 14799; with the last 3 digits like that, we thought it would
be a good one.
But, the TeVatron went down with a Central Helium Liquifier problem
and stayed off for the rest of the shift.

Description of Shift:

run spill time comment
14798 16:00 End run, tapes full.
14799 16:15 Start run. New tapes. UP* 64-66, UPS022.

94 18:15 R.K. and B.T. looked at HA2 on a scope and it's
scalar count suddenly droped by x2 and is now
back to it's original value (i.e., it had been
counting x2 high, but trigger and veto fractions
never changed. Problem most likely localized to

19:20 Sasha L. changed TRD HV so MIP peak is at 40 counts.
14799 199 20:35 End Run.

14800 20:55 Start Long Laser Scan.
14800 21:40 Laser scan ended.

14801 21:45 Start cosmics (no output) to keep us warm.
22:00 Access to NM2 to remove lead from Acc. Counters.
22:15-22:30 R.K tests out the new couch: "It's ok, but we need a
canopy to block the lights; or a dimmer switch."

Follow up Notes on Previous Shift

Items to Follow up on Next Shift